OpenStudy (miss_skeelo):

Can someone please help me with 9 questions? Which point is on the graph of y = -2x + 1? Question 2 options: (2, 7) (3, -8) (6, 8) (4, -7)

rishavraj (rishavraj):

try the options one by one...like in 1st option (2, 7) means x = 2 and y = 7 plug those values in the equation and check if it satisfies

OpenStudy (miss_skeelo):

I'm not sure how to do that though...

rishavraj (rishavraj):

lets check 1st option x = 2, y = 7 plug them in the equation y = -2x + 1 7 = -2(2) + 1 7 = -4 + 1 7 = -3 which is wrong....tht means 1st option is incorrect